The Data of Macroeconomics-End of Chapter Problem Abby consumes only apples. In year 1, red apples cost $1 each, green apples cost $2 each, and Abby buys 10 red apples. In year 2, red apples cost $2 each, green apples cost $1 each, and Abby buys 10 green apples. Assume that year 1 is the base year in which the consumer basket is fixed. a. Compute the CPI for apples for each year. Assume that year 1 is the base year in which the consumer basket is fixed. How does your index change from year 1 to year 2? The CPI for year is:1 The CPI for year 2 is: 2 CPl hasdoubled. b. Compute Abbys nominal spending on apples in cach year. How does it change from year 1 to year 2? Nominal Spending in year : $ 10 Nominal spending in year 2: 10 Nominal spending has stayed the same. c. Using year 1 as the base year, compute Abbys real spending on apples in each year. How does it change from year 1 to year 2? Real spending in year 1: S 10 Real spending in year 2:$ 20 Real spending hasdoubled.

d. Defining the implicit price deflator as nominal spending divided by real spending, compute the deflator for each year. How does the deflator change from year 1 to year 2? Deflator in year : 1 Deflator in year 2: 0.5 The deflator has fallen by half. e. Suppose that Abby is equally happy eating red or green apples. How much has the true cost of living increased for Abby? Compare this answer to your answers to parts (a) and (d). The true cost of living has stayed the same.

Need help with this, CORRECT ANSWERS ARE ALREADY SHOWN. All I need to know is how to arrive at these answers, please show work
Answer #1
Year Red apples cost Green apples cost Red apples bought Green apples bought Nominal spending Real spending Deflator
1 1 2 10 nil

10*1+2*0

= 10

10*1+2*0

=10

(10/10)=1
2 2 1 nil 10

2*0+10*1

= 10

0*1+10*2

=20

(10/20)=0.5

Nominal spending = Current price * Quantity in that year

Real spending = Base price * Quantity in that year

Base price = Price in year 1

Deflator = Nominal spending / Real spending

CPI = Cost of basket in current period/ Cost of basket in base period

Fix the basket for year 1 as 10 red apples

Cost of basket in year 1 = 10*1+2*0 = 10 (Base year)

Cost of basket in year 2 = 2*10+1*0= 20

CPI in year 1 = (10/10) = 1

CPI in year 2 = (20/10) = 2

It doubled.

Nominal GDP is same

Real GDP doubled

Deflator fallen by half

Cost of buying apples in both years is same for Abby.